# STI-Bench: Are MLLMs Ready for Precise Spatial-Temporal World Understanding?

### v3

@@ Abstract
The use of Multimodal Large Language Models (MLLMs) as an end-to-end solution for Embodied AI and Autonomous Driving has become a prevailing trend. While MLLMs have been extensively studied for visual semantic understanding tasks, their ability to perform precise and quantitative spatial-temporal understanding in real-world applications remains largely unexamined, leading to uncertain prospects. To evaluate models’ Spatial-Temporal Intelligence, we introduce STI-Bench, a benchmark designed to evaluate MLLMs’ spatial-temporal understanding through challenging tasks such as estimating and predicting the appearance, pose, displacement, and motion of objects. Our benchmark encompasses a wide range of robot and vehicle operations across desktop, indoor, and outdoor scenarios. The extensive experiments reveal that the state-of-the-art MLLMs still struggle in real-world spatial-temporal understanding, especially in tasks requiring precise distance estimation and motion analysis. Paper Page: https://mint-sjtu.github.io/STI-Bench.io/
@@ 1 Introduction
To answer this question, we propose a Spatial-Temporal Intelligence Benchmark (STI-Bench), designed to evaluate MLLMs’ spatial-temporal world understanding capability. We evaluate MLLMs using single video or multiple images as input instead of 3D point clouds. The main reasons are: 1) the majority of state-of-the-art models, e.g., GPT-4o [31] and Gemini [36], can accept images or video as input rather than 3D point clouds; 2) Videos are more frequently used in human’s daily life and they usually contain sufficient information to infer the spatial-temporal environment. STI-Bench contains 300 videos and more than 2,000 QA pairs, covering three major scenarios: Desktop, Indoor, and Outdoor. The videos are sourced from Omni6DPose [49], ScanNet [16] and Waymo [34] respectively, thus encompassing a broad spectrum of real-world environments. As illustrated in Figure 2, we design eight distinct tasks to evaluate models’ ability of static spatial measurement and grounding, and dynamic tasks including speed, acceleration and trajectory estimation.
@@
  • We present STI-Bench, comprising over 300 videos and more than 2,000 tailored questions across desktop, indoor, and outdoor scenarios, providing a systematic quantitative assessment of MLLMs’ spatial-temporal understanding capabilities.
@@ 2.1 Multimodal Large Language Models
Multimodal large language models (MLLMs) have achieved groundbreaking performance in visual understanding [1, 4, 12, 36], leveraging large language models (LLMs) [42, 37, 35] and visual encoders. Beyond image-based MLLMs, recent advancements have extended multimodal learning to video understanding. Classical works include models like VideoChat[23], which enable interactive video-based dialogue by integrating multimodal understanding. Subsequent models like Subsequent models like Video-LLaVA[28] enhance visual-language alignment through large-scale vision-language pretraining and fine-tuned adaptation, extending LLaVA[29]’s capability to process video inputs effectively. Recent works, Qwen2.5-VL [42] excels in long-video understanding and temporal localization by incorporating absolute temporal encoding, enabling the model to capture relationships among video frames more effectively. Additionally, its advancement in dynamic resolution modeling allows for seamless adaptation to videos with varying sampling rates, enhancing its versatility in processing diverse video inputs.
@@ 3.2 Benchmark Construction
To encompass a broad spectrum of real-world environments, STI-Bench covers three major scenarios: Desktop, Indoor, and Outdoor. Accordingly, we draw from three publicly available datasets—Omni6DPose[49] for desktop-scale 6D object pose estimation, ScanNet[16] for indoor 3D scene reconstruction, and Waymo [34] for autonomous driving. These datasets provide frame-by-frame camera intrinsic and extrinsic parameters, as well as point clouds for each object, which we map to two-dimensional bounding boxes in each frame.
@@
After generating and refining the QA pairs, we recognized that real-world applications differ significantly in terms of error tolerance. For instance, a desktop robotic arm may require millimeter-level precision, whereas autonomous driving can function effectively with meter-scale accuracy. To accommodate these varied needs, we applied a scaling factor to the numerical differences between correct answers and distractors, aligning them with the precision requirements of specific scenarios. Consequently, the smallest margin of error ranges from millimeters to centimeters in desktop settings, centimeters to decimeters indoors, and decimeters to meters outdoors. We also adopted a logarithmic sampling approach to avoid clustering most differences at the higher bounds of each range. This fine-grained adjustment preserves the semantic value of each question while maintaining suitable gradients across different precision levels, enabling more effective training and evaluation of MLLMs in diverse environments and industries.
@@
GPT-4o[31] 5 34.8 24.9 49.6 28.1 27.6 36.0 30.3 36.8 51.3
Claude-3.7-Sonnet[2] 3 39.4 31.8 49.0 36.3 29.0 36.9 27.0 41.0 62.7
Gemini-2.0-Flash[36] 4 38.7 33.7 50.0 33.7 32.7 34.4 15.1 48.7 62.4
Gemini-2.5-Pro[36] 2 40.9 34.2 53.4 32.3 32.4 34.3 44.9 52.0 58.4
@@
VideoLLaMA3-7B[47] 8 26.9 25.6 36.3 24.3 22.4 27.5 23.8 28.2 31.7
MiniCPM-V-2.6[45] 8 26.9 27.7 44.5 29.0 19.0 25.7 7.0 30.8 35.6
InternVL2.5-78B[13] 7 28.4 22.1 38.9 26.0 23.2 30.2 26.5 32.4 36.1
VideoChat-Flash[26] 6 34.1 27.9 21.2 29.7 27.4 22.3 64.7 45.8 46.4
Qwen2.5-VL-72B[5] 1 41.3 32.2 46.2 39.6 28.3 35.6 50.3 44.9 61.0
Table 2: Evaluation on Ourbench. Orange marks the best result, and Light Orange marks the second best.
@@
VideoLLaMA3-7B 26.61 27.63 25.86 26.89
MiniCPM-V-2.6 27.87 26.13 26.60 26.89
InternVL2.5 34.11 25.89 23.81 28.39
VideoChat-Flash 36.12 35.32 27.53 34.08
GPT-4o 39.12 34.45 27.41 34.82
Claude-3.7-Sonnet 43.43 36.53 37.41 39.35
Gemini-2.0-Flash 40.71 37.57 37.41 38.74
Gemini-2.5-Pro 47.48 38.55 33.09 40.88
Qwen2.5-VL 49.24 35.63 36.23 41.32
@@ 4.1 Settings
We conduct a thorough evaluation of leading MLLMs from diverse model families, focusing on both proprietary and open-source solutions. Specifically, we assess the performance of four proprietary models, GPT-4o[31], Gemini-2.0-Flash[36], Gemini-2.5-Pro[36], and Claude-3.7-Sonnet[2], as well as several representative open-source MLLMs that have undergone specialized video-related training, including Qwen2.5-VL-72B[5], InternVL2.5-78B[13], MiniCPM-V-2.6[45], VideoChat-Flash[26] and VideoLLaMA3-7B[47].
Considering the stability of open-source models and the API limitations of proprietary models, we uniformly sample 30 frames from the video for each record and explicitly indicate the sampling FPS (Frames Per Second) for the current video within the prompt. An exception is made for Claude-3.7-Sonnet, for which only 20 frames are sampled due to its API constraints. Our benchmark tasks are presented in a multiple-choice format with five possible answers, hence a random guess baseline yields a 20% accuracy. We measure each model’s accuracy by directly comparing the model’s selected answer with the ground truth, without employing any additional external models or annotations for performance evaluation.
@@ 4.2 Main Results
As shown in Table 2, we present a comprehensive evaluation of various MLLMs on on STI-Bench. Overall, Qwen2.5-VL-72B achieves the highest average accuracy of 41.3% among all tested models, slightly outperforming Gemini-2.5-Pro (40.9%). While these results significantly exceed the random guess baseline (20% ), they still highlight substantial room for improvement in spatial-temporal understanding.
When analyzing performance across different scene types, we observe varied patterns. Qwen2.5-VL demonstrates significantly better performance in outdoor scenarios (49.2%) compared to its indoor (35.6%) and desktop performance (36.2%). Conversely, Gemini-2.5-Pro shows its strongest performance in indoor environments (38.6%), followed by outdoor (47.5%) and then desktop scenes (33.1%). Other models like Claude-3.7-Sonnet and Gemini-2.0-Flash show strength in desktop scenarios (both 37.4%). This suggests varying model specializations, possibly influenced by training data distribution or architectural differences in handling different environmental cues.
Task-specific performance reveals particularly challenging areas. Models generally struggle with tasks requiring precise quantitative estimation. For instance, the best performance on Displacement & Path Length is achieved by Gemini-2.0-Flash (32.7%), and on Dimensional Measurement by Gemini-2.5-Pro (34.2%). Speed & Acceleration also presents difficulties, with Claude-3.7-Sonnet leading at 36.9%. In contrast, models demonstrate stronger capabilities in Pose Estimation (best: 62.7% by Claude-3.7-Sonnet) and Spatial Relation tasks (best: 53.4% by Gemini-2.5-Pro). Notably, VideoChat-Flash achieves the highest score in Ego Orientation (64.7%) despite its lower overall average.
Among open-source models, Qwen2.5-VL stands out, demonstrating highly competitive performance and achieving the top overall rank. However, other open-source models like InternVL2.5, VideoLLaMA3-7B, and MiniCPM-V-2.6 still lag significantly behind the top performers and larger counterparts.
It is important to emphasize that even the best-performing model, Qwen2.5-VL, achieves only 41.3% average accuracy on our benchmark. While this is considerably better than random guessing, it remains far from the reliability required for demanding real-world applications like embodied AI or autonomous driving. These results indicate that current MLLMs, despite their impressive general visual understanding capabilities, still require significant advancements in precise spatial-temporal intelligence for embodied tasks.
@@ 4.3 Experimental Analysis
Given that Gemini-2.5-Pro exhibits strong multi-modal reasoning capabilities, provides detailed thinking processes (as exemplified in Figure 6), and ranks second-best overall (best among proprietary models), we select it as a representative for in-depth analysis.
Our analysis of Gemini-2.5-Pro reveals several key characteristics of current state-of-the-art MLLMs’ spatial-temporal understanding capabilities. Overall, the model achieves 40.9% average accuracy. Performance varies across scene types, with the strongest results in indoor scenarios (38.6%), followed by outdoor (47.5%) and desktop environments (33.1%). This performance disparity suggests that the model’s training data likely emphasized outdoor scenes and larger-scale understanding, or that outdoor environments typically provide clearer visual cues for spatial reasoning.
When examining task-specific performance for Gemini-2.5-Pro (Table 2), we observe its strongest capabilities in tasks involving relative understanding and object state. It achieves its highest score in Pose Estimation (58.4%), followed by Spatial Relation (53.4%, also the best overall score for this task among all models). Its performance in Ego Orientation is 44.9%. Interestingly, while tasks requiring precise quantitative estimation generally show lower absolute scores across models, Gemini-2.5-Pro performs very competitively or even leads in these areas: it achieves the top score in Dimensional Measurement (34.2%) and the second-best score in Displacement & Path Length (32.4%). Its score for Speed & Acceleration is 34.3%. This indicates that while the absolute accuracy in precise estimation tasks remains a challenge, Gemini-2.5-Pro shows comparatively strong capabilities in these domains relative to other models.
@@
These error patterns highlight that, despite Gemini-2.5-Pro’s strong performance relative to other models, it still faces significant challenges in precise spatial-temporal understanding. It’s limitations in quantitative estimation and complex spatial-temporal reasoning indicate that current MLLMs remain far from achieving the reliability required for embodied AI or autonomous driving applications.
@@ 5 Conclusion
We introduced STI-Bench, a comprehensive benchmark to assess MLLMs’ spatial-temporal understanding through over 300 real-world videos and 2,000 QA pairs of robot desktop, indoor, and outdoor scenarios, which reveals significant limitations in current MLLMs’ spatial-temporal understanding capabilities, with even the top-performing models achieving no more than 50% accuracy. Models particularly struggle with precise quantitative tasks like dimensional measurement. Our analysis identifies three key weaknesses: inaccurate spatial quantification, flawed temporal dynamics understanding, and weak cross-modal integration. These findings emphasize the substantial gap between current capabilities and the reliability needed for embodied AI and autonomous driving applications. STI-Bench provides a valuable framework for evaluating and improving MLLMs’ ability to understand the physical world—essential for developing the next generation of embodied intelligent systems.

### v4

@@ Abstract
The use of Multimodal Large Language Models (MLLMs) as an end-to-end solution for Embodied AI and Autonomous Driving has become a prevailing trend. While MLLMs have been extensively studied for visual semantic understanding tasks, their ability to perform precise and quantitative spatial-temporal understanding in real-world applications remains largely unexamined, leading to uncertain prospects. To evaluate models’ Spatial-Temporal Intelligence, we introduce STI-Bench, a benchmark designed to evaluate MLLMs’ spatial-temporal understanding through challenging tasks such as estimating and predicting the appearance, pose, displacement, and motion of objects. Our benchmark encompasses a wide range of robot and vehicle operations across outdoor, indoor, and desktop scenarios. The extensive experiments reveal that the state-of-the-art MLLMs still struggle in real-world spatial-temporal understanding, especially in tasks requiring precise distance estimation and motion analysis. Paper Page: https://mint-sjtu.github.io/STI-Bench.io/
@@ 1 Introduction
To answer this question, we propose a Spatial-Temporal Intelligence Benchmark (STI-Bench), designed to evaluate MLLMs’ spatial-temporal world understanding capability. We evaluate MLLMs using single video or multiple images as input instead of 3D point clouds. The main reasons are: 1) the majority of state-of-the-art models, e.g., GPT-4o [33] and Gemini [38], can accept images or video as input rather than 3D point clouds; 2) Videos are more frequently used in human’s daily life and they usually contain sufficient information to infer the spatial-temporal environment. STI-Bench contains 300 videos and more than 2,000 QA pairs, covering three major scenarios: Desktop, Indoor, and Outdoor. The videos are sourced from Waymo [36], ScanNet [16], and Omni6DPose [51] respectively, thus encompassing a broad spectrum of real-world environments. As illustrated in Figure 2, we design eight distinct tasks to evaluate models’ ability of static spatial measurement and grounding, and dynamic tasks including speed, acceleration and trajectory estimation.
@@
  • We present STI-Bench, comprising over 300 videos and more than 2,000 tailored questions across outdoor, indoor, and desktop scenarios, providing a systematic quantitative assessment of MLLMs’ spatial-temporal understanding capabilities.
@@ 2.1 Multimodal Large Language Models
MLLMs have achieved groundbreaking performance in visual understanding [1, 4, 12, 38], leveraging large language models (LLMs) [44, 39, 37] and visual encoders. Beyond image-based MLLMs, recent advancements have extended multimodal learning to video understanding. Classical works include models like VideoChat[24], which enable interactive video-based dialogue by integrating multimodal understanding. Subsequent models like Subsequent models like Video-LLaVA[30] enhance visual-language alignment through large-scale vision-language pretraining and fine-tuned adaptation, extending LLaVA[31]’s capability to process video inputs effectively. Recent works, Qwen2.5-VL [44] excels in long-video understanding and temporal localization by incorporating absolute temporal encoding, enabling the model to capture relationships among video frames more effectively. Additionally, its advancement in dynamic resolution modeling allows for seamless adaptation to videos with varying sampling rates, enhancing its versatility in processing diverse video inputs.
@@ 3.2 Benchmark Construction
To encompass a broad spectrum of real-world environments, STI-Bench covers three major scenarios: Desktop, Indoor, and Outdoor. Accordingly, we draw from three publicly available datasets—Waymo [36] for autonomous driving, ScanNet[16] for indoor 3D scene reconstruction, and Omni6DPose[51] for desktop-scale 6D object pose estimation. These datasets provide frame-by-frame camera intrinsic and extrinsic parameters, as well as point clouds for each object, which we map to two-dimensional bounding boxes in each frame.
@@
To ensure task diversity and challenge, each question is equipped with carefully constructed distractor options that are both significantly different from the correct answer while remaining within a reasonable range. Considering the varying precision requirements across different scenarios—millimeter-level for desktop applications, centimeter-to-decimeter-level for indoor environments, and decimeter-to-meter-level for outdoor scenes—we applied scenario-specific scaling factors to generate distractors and employed logarithmic sampling techniques to distribute numerical differences more evenly, thereby enhancing the robustness of our evaluation.
Specifically, given a scene type $S\in\{\text{Desktop},\text{Indoor},\text{Outdoor}\}$, a correct answer $A_{correct}$, and four initial distractor options $\{A_{d1},A_{d2},A_{d3},A_{d4}\}$, our fine-grained adjustment process consists of the following steps:
[…]
2. We logarithmically sample an error value $e$ from the corresponding error range:
$e=E_{min}\cdot(E_{max}/E_{min})^{u}$
where $u\sim\mathcal{U}(0,1)$ is a uniform random number.
3. We adjust each distractor option $A_{di}$ to $A_{di}^{\prime}$ using a weighted average:
$A_{di}^{\prime}=(1-w)\cdot A_{di}+w\cdot A_{correct}$
where the weight $w$ is identical for all distractors and is determined by solving:
$\min_{i\in\{1,2,3,4\}}\|A_{di}^{\prime}-A_{correct}\|=e$
This yields:
$w=1-\frac{e}{\min_{i\in\{1,2,3,4\}}\|A_{di}-A_{correct}\|}$
This approach ensures that the minimum distance between any adjusted distractor and the correct answer equals the sampled error value $e$, while other distractors maintain relatively larger distances but are adjusted using the same weight factor.
@@ 3. We adjust each distractor option $A_{di}$ to $A_{di}^{\prime}$ using a weighted average:
GPT-4o[33] 8 34.8 27.1 51.8 29.0 23.2 35.4 33.7 32.0 53.6
Gemini-2.0-Flash[38] 4 38.7 31.9 50.0 31.8 27.7 32.1 10.8 38.5 61.3
Claude-3.7-Sonnet[2] 3 40.5 29.8 45.5 35.7 28.9 38.8 40.0 47.4 62.6
Gemini-2.5-Pro[20] 1 41.4 38.7 53.8 36.9 33.9 33.1 52.5 47.4 50.4
@@
MiniCPM-V-2.6[47] 10 26.9 27.7 44.5 29.0 19.0 25.7 7.0 30.8 35.6
VideoChat-R1[28] 9 32.8 23.2 47.3 31.5 22.4 31.1 26.0 47.9 48.3
VideoLLaMA3-7B[49] 7 35.2 29.4 48.6 36.1 21.5 36.7 23.2 54.6 48.1
VideoChat-Flash[27] 6 36.3 33.6 51.4 33.1 27.1 32.3 22.2 54.2 51.4
InternVL2.5-78B[13] 5 38.5 29.9 52.8 31.6 24.9 37.2 49.2 43.6 53.6
Qwen2.5-VL-72B[5] 2 40.7 31.5 47.6 39.1 25.1 38.4 43.8 51.3 60.6
Table 2: Evaluation on STI-Bench. Orange = best; Light Orange = second best.Note: The random baseline for all tasks is 20%.
@@
MiniCPM-V-2.6[47] 27.9 26.1 26.6 26.9
VideoChat-R1_7B[28] 41.1 28.0 27.1 32.8
GPT-4o[33] 41.4 33.1 27.3 35.2
VideoLLaMA3-7B[49] 41.2 31.9 30.5 35.2
Gemini-2.0-Flash[38] 39.1 35.9 30.5 36.0
VideoChat-Flash[27] 38.6 36.7 31.3 36.3
InternVL2.5-78B[13] 46.1 34.5 32.1 38.5
Claude-3.7-Sonnet[2] 47.2 38.2 32.3 40.5
Qwen2.5-VL-72B[5] 50.6 35.0 33.5 40.7
Gemini-2.5-Pro[20] 48.7 37.1 35.8 41.4
@@ 4.1 Settings
We conduct a thorough evaluation of leading MLLMs from diverse model families, focusing on both proprietary and open-source solutions. Specifically, we assess the performance of four proprietary models, GPT-4o[33], Gemini-2.0-Flash[38], Gemini-2.5-Pro[38], and Claude-3.7-Sonnet[2], as well as several representative open-source MLLMs that have undergone specialized video-related training, including Qwen2.5-VL-72B[5], InternVL2.5-78B[13], MiniCPM-V-2.6[47], VideoChat-Flash[27],VideoChat-R1[28] and VideoLLaMA3-7B[49].
Considering the stability of open-source models and the API limitations of proprietary models, we uniformly sample 30 frames from the video for each record and explicitly indicate the sampling FPS for the current video within the prompt. An exception is made for Claude-3.7-Sonnet, for which only 20 frames are sampled due to its API constraints. Our benchmark tasks are presented in a multiple-choice format with five possible answers, hence a random guess baseline yields a 20% accuracy. We measure each model’s accuracy by directly comparing the model’s selected answer with the ground truth, without employing any additional external models or annotations for performance evaluation.
@@ 4.2 Main Results
As shown in Table 2 and Table 3, we present a comprehensive evaluation of various MLLMs on STI-Bench. Overall, Gemini-2.5-Pro achieves the highest average accuracy of 41.4% among all tested models, closely followed by Qwen2.5-VL-72B at 40.7% among open-source models. While these results significantly exceed the random guess baseline (20%), they still highlight substantial room for improvement in spatial-temporal understanding.
When analyzing performance across different scene types (Table 3), we observe varied patterns. Qwen2.5-VL-72B demonstrates notably strong performance in outdoor scenarios (50.6%), while Gemini-2.5-Pro also performs well outdoors (48.7%). In indoor environments, Claude-3.7-Sonnet (38.2%) and Gemini-2.5-Pro (37.1%) lead. For desktop scenes, Gemini-2.5-Pro shows the strongest performance (35.8%), followed by Qwen2.5-VL-72B (33.5%). These variations suggest different model specializations, possibly influenced by training data distribution or architectural differences in handling diverse environmental cues.
Task-specific performance (Table 2) reveals particularly challenging areas. Models generally struggle with tasks requiring precise quantitative estimation. For instance, the best performance on Dimensional Measurement is achieved by Gemini-2.5-Pro (38.7%), and on Displacement & Path Length also by Gemini-2.5-Pro (33.9%). Speed & Acceleration also presents difficulties, with Claude-3.7-Sonnet leading at 38.8%. In contrast, models demonstrate stronger capabilities in Pose Estimation (best: 62.7% by Claude-3.7-Sonnet) and Spatial Relation tasks (best: 53.8% by Gemini-2.5-Pro). Notably, Gemini-2.5-Pro also achieves the highest score in Ego-Centric Orientation (52.5%) among proprietary models, while VideoLLaMA3-7B leads in Trajectory Description (54.6%) among open-source models.
Among open-source models, Qwen2.5-VL-72B stands out with an average accuracy of 40.7%, demonstrating highly competitive performance. However, other open-source models like InternVL2.5-78B (38.5%) also show promise, though smaller models like MiniCPM-V-2.6 (26.9%) still lag behind their larger counterparts and proprietary models.
It is important to emphasize that even the best-performing model, Gemini-2.5-Pro, achieves only 41.4% average accuracy on our benchmark. While this is considerably better than random guessing, it remains far from the reliability required for demanding real-world applications like embodied AI or autonomous driving. These results indicate that current MLLMs, despite their impressive general visual understanding capabilities, still require significant advancements in precise spatial-temporal intelligence for embodied tasks.
@@ 4.3 Experimental Analysis
Given that Gemini-2.5-Pro exhibits strong multi-modal reasoning capabilities, provides detailed thinking processes (as exemplified in Figure 6), and achieves the top overall performance (41.4% average accuracy) among all evaluated models, we select it as a representative for in-depth analysis.
Our analysis of Gemini-2.5-Pro reveals several key characteristics of current state-of-the-art MLLMs’ spatial-temporal understanding capabilities. Overall, the model achieves 41.4% average accuracy. Performance varies across scene types: it performs strongest in outdoor scenarios (48.7%), followed by indoor (37.1%) and desktop environments (35.8%). This performance disparity might suggest that the model’s training data or architecture is better attuned to outdoor and larger-scale understanding, or that such environments typically provide clearer visual cues for spatial reasoning.
When examining task-specific performance for Gemini-2.5-Pro (Table 2), we observe its strongest capabilities in tasks involving relative understanding and object state. It achieves high scores in Spatial Relation (53.8%, the best among all models) and Ego-Centric Orientation (52.5%, the highest among proprietary models), followed by Pose Estimation (50.4%). Interestingly, while tasks requiring precise quantitative estimation generally show lower absolute scores across models, Gemini-2.5-Pro performs very competitively and often leads in these areas: it achieves the top scores in Dimensional Measurement (38.7%) and Displacement & Path Length (33.9%). Its score for Speed & Acceleration is 33.1%. This indicates that while the absolute accuracy in precise estimation tasks remains a challenge, Gemini-2.5-Pro demonstrates comparatively strong capabilities in these demanding domains relative to other models.
@@
These error patterns highlight that, despite Gemini-2.5-Pro’s strong performance relative to other models, it still faces significant challenges in precise spatial-temporal understanding. Its limitations in quantitative estimation and complex spatial-temporal reasoning indicate that current MLLMs remain far from achieving the reliability required for embodied AI or autonomous driving applications.
@@ 5 Conclusion
We introduced STI-Bench, a comprehensive benchmark to assess MLLMs’ spatial-temporal understanding through over 300 real-world videos and 2,000 QA pairs of robot outdoor, indoor, and desktop scenarios, which reveals significant limitations in current MLLMs’ spatial-temporal understanding capabilities, with top-performing models like Gemini-2.5-Pro achieving around 41.4% average accuracy. Models particularly struggle with precise quantitative tasks like dimensional measurement. Our analysis identifies three key weaknesses: inaccurate spatial quantification, flawed temporal dynamics understanding, and weak cross-modal integration. These findings emphasize the substantial gap between current capabilities and the reliability needed for embodied AI and autonomous driving applications. STI-Bench provides a valuable framework for evaluating and improving MLLMs’ ability to understand the physical world—essential for developing the next generation of embodied intelligent systems.